OMG Inc. has 4 million shares of common stock outstanding, 3 million shares of preferred stock outstanding, and 5,000 bonds. Suppose the common shares sell for $19 per share, the preferred shares sell for $18 per share, and the bonds sell for 108 percent of par.
What weight should you use for preferred stock in the computation of OMG’s WACC? (Round your answer to 2 decimal places.)
Weight used: ____.__%
Answer #1

Market value of common stock = 4,000,000 * 19 = 76,000,000

Market value of preferred stock = 3,000,000 * 18 = 54,000,000

Assuming face to be $1,000

Price = 108% of 1000 = 1080

Market value of debt = 5,000 * 1,080 = 5,400,000

Total market value = 76,000,000 + 54,000,000 + 5,400,000 = 135,400,000

Weight of preferred stock = (54,000,000 / 135,400,000) * 100

Weight of preferred stock = 39.88%